# Chern insulators and topological flat bands in cavity-embedded kagome systems

### v1

@@ V Conclusion and Outlook
We have studied the band topology of a kagome system using the muffin-tin potential under a circularly polarized cavity field. Because the cavity field, which breaks time-reversal symmetry, induces a mass gap, a Chern insulating phase emerges when the light-matter interaction is introduced. We have shown that a flat band acquires a nonzero Chern number due to the light-matter interaction. Moreover, we have found that stronger light-matter coupling can induce topological phase transitions between different Chern insulating phases in the kagome system. We also have demonstrated existence of topological edge states by constructing tight-binding models. The chirality of the edge states is switched when the topological phase transition occurs in the USC regime. Our work has revealed that various Chern insulating phases can be realized in the cavity-embedded kagome system.
We have investigated topological band structures in a cavity-embedded kagome system with a narrow bandwidth, where three low-energy bands are isolated from other energy bands in the presence of light-matter interaction. In the THz regime, such band structures may be realizable in heavy-fermion kagome materials with narrow bandwidths [60, 61]. In addition, Dirac points in kagome materials, such as CsV3Sb5 [62, 63, 64], have been experimentally observed. We expect that typical kagome systems with Dirac points, where the bandwidth is not sufficiently narrow, exhibit topological bands with nonzero Chern numbers in a circularly polarized cavity since the light-matter interaction opens a band gap due to time-reversal symmetry breaking. On the other hand, in such kagome materials, the photon vacuum can strongly hybridize with higher-energy $n$-photon states ($n>0$) in the (sub-)THz regime. This hybridization may induce topological phase transitions that lead to various topological band structures beyond those considered in this paper. A detailed analysis of the hybridization effect is left for future work. In addition, quantum Hall transport has been experimentally measured in split-ring resonator cavities [11, 13]. Therefore, the kagome Chern insulators induced by cavity fields can be observed via quantum Hall transport.
@@ A.1 Construction of the low-energy Hamiltonian
Before detailing the construction of the low-energy Hamiltonian in the kagome and honeycomb systems, we briefly explain the reason for using the AD frame [55, 28]. When $g/\omega_{c}\gtrsim 0.1$, the Peierls substitution may not accurately produce energy bands of the cavity QED Hamiltonian [55, 28]. Therefore, in such an USC regime, we need a different approach to construct a low-energy tight-binding model. In the AD frame, $\xi_{g}$ vanishes both in the weak-coupling limit $(g/\omega_{c}\rightarrow 0)$ and in the strong-coupling limit $(g/\omega_{\mathrm{c}}\rightarrow\infty)$, as seen in Eq. (6). Because electronic and photonic degrees of freedom are asymptotically decoupled in these limits, low-energy eigenstates can be approximated by the product state of an electronic wavefunction and the photon vacuum. Hence, the effective Hamiltonian in Eq. (8) can reproduce the low-energy band structure in the strong-coupling limit, which is useful to construct the tight-binding model.
@@
where $\bm{G}=m_{1}\bm{b}_{1}+m_{2}\bm{b}_{2}$ is a reciprocal lattice vector with integers $m_{1}$ and $m_{2}$ and $\bm{b}_{1}=2\pi/a(1,1/\sqrt{3})$ and $\bm{b}_{2}=2\pi/a(0,2/\sqrt{3})$. Here,
@@ A.2 Band topology of the tight-binding model
Thus, when $\xi_{g}/a$ is vanishingly small, the hybridization between the photon vacuum and the $n$-photon states becomes negligible, and does not significantly affect the band topology. According to Eq. (6), the effective coupling strength decreases as the frequency $\omega_{c}$ is higher. In our frequency regime, this hybridization does not change the possible topological phases. Therefore, because $\hat{H}^{\mathrm{AD}}_{0}$ can describe the low-energy band topology of our kagome system, the tight-binding model can reproduce its topological phases.
@@ Appendix B Edge states in the honeycomb Chern insulator
Figure 6: (a) Band structure in the tight-binding model for the honeycomb system at $g/\omega_{\mathrm{c}}=0.5$. We shift the lowest energy eigenvalue at the K point to match that obtained from the exact analysis. (b) Topological edge states with the periodic boundary condition in the $x$-direction and open boundary condition in the $y$-direction. The red (green) line indicates gapless edge modes localized at the top (bottom) edge.

### v2

@@ V Conclusion and Outlook
We have studied the band topology of a kagome system using the muffin-tin potential in a chiral cavity. Because the cavity field, which breaks time-reversal symmetry, induces a mass gap, a Chern insulating phase emerges when the light-matter interaction is introduced. We have shown that a flat band acquires a nonzero Chern number due to the light-matter interaction. Moreover, we have found that stronger light-matter coupling can induce topological phase transitions between different Chern insulating phases in the kagome system. These changes in the band topology stem from gap closings between different pairs of the three bands, suggesting that the presence of multiple bands is important. We also have demonstrated existence of topological edge states by constructing tight-binding models. The chirality of the edge states is switched when the topological phase transition occurs in the USC regime. Our work has revealed that various Chern insulating phases can be realized in the cavity-embedded kagome system. Experimentally, quantum Hall transport has been measured in split-ring resonator cavities [12, 14]. Therefore, the kagome Chern insulators induced by cavity fields can be observed via quantum Hall transport.
We have investigated topological band structures in a cavity-embedded kagome system. For simplicity, we have assumed that three low-energy bands with a narrow bandwidth are isolated from other energy bands in the presence of light-matter interaction. Recently, kagome-like band dispersions with a narrow bandwidth have been experimentally engineered in 2DEGs with patterned electrostatic potentials [65]. Moreover, chiral cavities have been experimentally realized [13, 71, 72, 73]. Therefore, in the (sub-)THz regime, 2DEGs embedded in chiral cavities can be regarded as a promising platform for realizing such cavity-induced topological phases  [^2] The vector potential amplitude $A_{0}$ is estimated as $A_{0}=\sqrt{\hbar/(2\epsilon_{0}\epsilon_{\mathrm{r}}\omega_{\mathrm{c}}V_{\mathrm{eff}})}$, where $\epsilon_{0}$ is the vacuum permittivity, $\epsilon_{\mathrm{r}}$ is the relative permittivity, and $V_{\mathrm{eff}}$ is the effective cavity mode volume. For InSb with $m\sim 0.01m_{\mathrm{e}}$ and $\epsilon_{\mathrm{r}}=15.7$, we can obtain the estimate $g/\omega_{\mathrm{c}}\sim 0.1$ for $V_{\mathrm{eff}}\sim 10^{-18}\ \mathrm{m}^{3}$ in the sub-THz regime. This suggests that the USC regime may be approached in cavities with a small effective mode volume, such as a metamaterial cavity with dipole nanoantennas [13] and the Faraday-rotator-based cavity [60]. .
For realistic materials, the above assumption is generally not satisfied in typical kagome materials  [^3] To satisfy the conditions in realistic materials with lattice constant $a\simeq 5$~Å in the THz regime, an effective mass ratio $m/m_{\mathrm{e}}$ of order $10^{2}$ is required in our model calculations. , although narrow bandwidths may be realized in heavy-fermion kagome materials [76, 77]. Nevertheless, we expect that typical kagome systems with Dirac points exhibit topological bands with nonzero Chern numbers in a chiral cavity since the light-matter interaction opens a band gap due to time-reversal symmetry breaking. Indeed, Dirac points have been experimentally observed in kagome materials, such as CsV3Sb5 [78, 79, 80]. Moreover, realistic kagome materials possess multiple bands near the Dirac points and flat band. Based on our results, the presence of multiple bands is expected to enrich the band topology through gap closings between different band pairs.
Finally, we comment on several effects that may render topological phase transitions accessible at moderate coupling strength. In realistic kagome materials, the photon vacuum can strongly hybridize with higher-energy $n$-photon sectors ($n>0$) in the (sub-)THz regime. This hybridization may induce topological phase transitions that lead to various topological band structures beyond those considered in this paper. Another possible route is to include multiple cavity modes. Because the effective mass $m_{\mathrm{eff}}$ can be enhanced by multiple modes [29, 81], this multimode effect can modify the band structures and may enable different topological phase transitions. A detailed analysis of these hybridization and multimode effects is left for future work.
@@ A.1 Construction of the low-energy Hamiltonian
Before detailing the construction of the low-energy Hamiltonian in the kagome and honeycomb systems, we briefly explain the reason for using the AD frame [61, 29]. When $g/\omega_{\mathrm{c}}\gtrsim 0.1$, the Peierls substitution may not accurately produce energy bands of the cavity QED Hamiltonian [61, 29]. Therefore, in such an USC regime, we need a different approach to construct a low-energy tight-binding model. In the AD frame, $\xi_{g}$ vanishes both in the weak-coupling limit $(g/\omega_{\mathrm{c}}\rightarrow 0)$ and in the strong-coupling limit $(g/\omega_{\mathrm{c}}\rightarrow\infty)$, as seen in Eq. (6). Because electronic and photonic degrees of freedom are asymptotically decoupled in these limits, low-energy eigenstates can be approximated by the product state of an electronic wavefunction and the photon vacuum. Hence, the effective Hamiltonian in Eq. (8) can reproduce the low-energy band structure in the strong-coupling limit, which is useful to construct the tight-binding model.
@@
where $\bm{G}=m_{1}\bm{b}_{1}+m_{2}\bm{b}_{2}=(G_{x},G_{y})$ is a reciprocal lattice vector with integers $m_{1}$ and $m_{2}$ and $\bm{b}_{1}=2\pi/a(1,1/\sqrt{3})$ and $\bm{b}_{2}=2\pi/a(0,2/\sqrt{3})$. Here,
@@ A.2 Band topology of the tight-binding model
Thus, when $\xi_{g}/a$ is vanishingly small, the hybridization between the photon vacuum and the $n$-photon sectors becomes negligible, and does not significantly affect the band topology. According to Eq. (6), the effective coupling strength decreases as the frequency $\omega_{\mathrm{c}}$ is higher. In our frequency regime, this hybridization does not change the possible topological phases. Therefore, because $\hat{H}^{\mathrm{AD}}_{0}$ can describe the low-energy band topology of our kagome system, the tight-binding model can reproduce its topological phases.
Figure 7: (a) Band structure in the tight-binding model for the honeycomb system at $g/\omega_{\mathrm{c}}=0.5$. We shift the lowest energy eigenvalue at the $\mathrm{K}$ point to match that obtained from the exact analysis. (b) Topological edge states with the periodic boundary condition in the $x$-direction and open boundary condition in the $y$-direction. Red and green solid lines represent gapless edge modes, and the color difference indicates that these modes are localized at opposite edges.